A total of $3000 is invested, part at 2% simple interest and part at 4%. If the total
annual return from the two investments is $100, how much is invested at each
rate?

Answer by Earlsdon(6294)

Let x = the amount invested at 2% while the remainder ($3,000-x) is the amount invested at 4%.
The interest earned on the two amounts can be expressed as:
0.02x and 0.04(3000-x) and the sum of these earnings is $100.00
Here's the equation you'll need to solve for x.
0.02x+0.04(3000-x) = 100 simpify.
0.02x+120-0.04x = 100 Combine the x-terms.
-0.02x+120 = 100 Subtract 120 from both sides.
-0.02x = -20 Finally, divide both sides by -0.02
x = 1000 and 3000-x = 2000
$1000 was invested at 2% and $2000 was invested at 4%